OpenStudy (anonymous):

Which of these frequency counts would be represented by the tallest bar on a histogram? A. 20 B. 21 C. 22 D. 19

OpenStudy (daniellelovee):

biggest number

OpenStudy (daniellelovee):

was it right or was it left? :)

OpenStudy (anonymous):

22 is the biggest number...

OpenStudy (daniellelovee):

yup

OpenStudy (anonymous):

If the data below were represented with a stem-and-leaf plot, a stem of 3 would have how many leaves? 53, 34, 16, 93, 72, 30, 23, 48, 31, 83 A. 5 B. 4 C. 2 D. 3

OpenStudy (anonymous):

@Daniellelovee

OpenStudy (daniellelovee):

all numbers divisible by 3

OpenStudy (anonymous):

D?

OpenStudy (daniellelovee):

lol divisible not number 3 XD when I say divisible I mean that it can be divided by 3

OpenStudy (daniellelovee):

there are 4

OpenStudy (daniellelovee):

and thank you for the medal :)

OpenStudy (anonymous):

A is it A

OpenStudy (daniellelovee):

o.O there are only 4 numbers that can be divided by 3 how did you get 5?

OpenStudy (anonymous):

psh i guessed.

OpenStudy (anonymous):

half my answers i have guessed. lol

OpenStudy (anonymous):

on other things.. xD

OpenStudy (daniellelovee):

lol well at least your honest XD

OpenStudy (daniellelovee):

the answer is B :p

OpenStudy (anonymous):

nope.. 3

OpenStudy (daniellelovee):

o.O whaaat???

OpenStudy (anonymous):

the answer was 3

OpenStudy (daniellelovee):

53=53/3 34=34/3 16=16/3 93=31--- 72=24--- 30=10--- 23 =23/3 48=16--- 31=31/3 83=83/3 and Im sorry but I object to that answer lol

OpenStudy (daniellelovee):

sorry :(

OpenStudy (anonymous):

Which of these is an example of discrete data? A. Area of a room B. Rides in an amusement park C. Temperature D. Distance

OpenStudy (anonymous):

@Daniellelovee

OpenStudy (daniellelovee):

I think C but idk

OpenStudy (anonymous):

B

OpenStudy (daniellelovee):

o.O ok I'm getting pissed that is 2 wrong in a row can you still get an A?

OpenStudy (anonymous):

i got 3 wrong so a C

OpenStudy (daniellelovee):

-.-' ugh I don't like that lol I'm a straight A's student! do you have any more quizzes?

OpenStudy (anonymous):

whole lesson

OpenStudy (daniellelovee):

alright go for it then

OpenStudy (anonymous):

kk i got a low c on it upper D. So at least i passed it.

OpenStudy (daniellelovee):

true but I don't like that grade lol lets try to fix it

OpenStudy (anonymous):

were not fixing that quix no no no were moving on

OpenStudy (daniellelovee):

haha ok lol

OpenStudy (anonymous):

sorry i just seen your comment

OpenStudy (daniellelovee):

is ok :)

OpenStudy (anonymous):

ok :)